OpenStudy (anonymous):

Solve the system algebraically. 2x+ y - 10 = 0 x - y - 4 = 0 What is the value of y? 1/3 2/3

OpenStudy (anonymous):

Use the method of elimination to solve this system of simultaneous equations.

OpenStudy (anonymous):

Adding equation 1 and equation 2 gives, 3x - 14 = 0. So x = 14/3, then plugging x into equation 2, you get y = x - 4 = 14/3 - 4 = 2/3

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

Generally, for these types of questions, you have two algebraic methods to solve them. One is by elimination, and the other is by substitution. Using elimination, you aim to add or subtract the equations to eliminate one of the variables to make the system single variable.